1.Engineers must consider the breadths of male heads when designing helmets. The company researchers have determined that the population of potential clientele have head breadths that are normally distributed with a mean of 7-in and a standard deviation of 1.2-in. Due to financial constraints, the helmets will be designed to fit all men except those with head breadths that are in the smallest 1.2% or largest 1.2%. What is the minimum head breadth that will fit the clientele? min = What is the maximum head breadth that will fit the clientele? max = Enter your answer as a number accurate to 1 decimal place.

2.Company XYZ know that replacement times for the quartz time pieces it produces are normally distributed with a mean of 16.9 years and a standard deviation of 1.8 years.

Find the probability that a randomly selected quartz time piece will have a replacement time less than 11.9 years?
P(X < 11.9 years) =

Enter your answer accurate to 4 decimal places.

If the company wants to provide a warranty so that only 2.5% of the quartz time pieces will be replaced before the warranty expires, what is the time length of the warranty?
warranty =  years

Enter your answer as a number accurate to 1 decimal place.
Answer #1

1)

for 1.2 % end values , critical z =-2.26

minimum head breadth that will fit the clientele =7-2.26*1.2 =4.3

maximum head breadth that will fit the clientele =7+2.26*1.2 =9.7

2)

a)

for normal distribution z score =(X-μ)/σx
here mean=       μ= 16.9
std deviation   =σ= 1.8
probability =P(X<11.9)=(Z<(11.9-16.9)/1.8)=P(Z<-2.78)=0.0027

b)

for 5th percentile critical value of z= -1.645
therefore corresponding value=mean+z*std deviation= 13.9